An aqueous solution of HCl is added to an aqueous solution of Na(OH) and a reaction occurs. The

concentration of H + ions in the solution is 1.0x10 -5 M, calculate the concentration of the OH –1 ions. Is the

solution more acidic or basic?


1
Expert's answer
2022-06-10T16:04:03-0400

pH = -log [H+]

pOH = -log [OH-]

in aqueous solution pH + pOH = 14


Using the above data, we find first the pH, then the pOH and then the OH-concentration.

pH = -log (1*10-5) = 5


pOH = 14 - pH = 14 - 5 = 9


pOH = -log [OH-] = 9

[OH-] = 1 * 10-9 M

If the pH is higher than 7, the solution is alkaline; if it is less than 7, it is acidic; if it is equal to 7, it is neutral.


Answer:

[OH-] = 1 * 10-9

solution is a acidic.